# Unveiling the Influence of Amplifying Language-Specific Neurons

### v2

@@ 5.1.3 Perplexity Changes
Self-intervention of LAPE neurons using pmax can modestly reduce perplexity in most languages, showing success in improving language modeling performance as illustrated in Figure 7. Perplexity tends to increase more in cross-intervention settings than in self-intervention. However, the extent of the increase is not always related to the similarity of the language pair; For example, intervention between highly similar same-family languages like id and ms does not lower perplexity, while less-related pairs like fr on en can. Interestingly, smaller models exhibit larger increases in perplexity, a pattern consistent with findings from the translation task intervention. This may suggest that smaller models contain more localized neurons responsible for language identity (see Appendix F for additional results).
@@ 5.2 Overlap Patterns in Language-Specific Neurons
As illustrated in (Figure 8), although LAPE neurons are inherently tailored to individual languages, overlaps still exist across languages, albeit at significantly lower proportions. Compared with Baseline neurons with higher overlap proportions (Appendix D), the stronger specificity of LAPE neurons significantly improve self-intervention and capture patterns in cross-intervention for perplexity and task performance, indicating that better specificity leads to better performance.
A strong overlap is captured in LAPE neurons across mutually intelligible languages like Indonesian (id) and Malaysian (ms), likely due to their linguistic similarity Ahmad and Pustaka (2011). Languages sharing writing systems, such as Chinese (zh) and Japanese (jp) Taylor et al. (1995), also show overlap, consistent with Kojima et al. (2024). Romance languages, i.e., Italian (it), French (fr), and Portuguese (pt), share neurons, and Vietnamese (vi) overlaps moderately with Thai (th) and possibly id, ms, jp, and zh, likely due to regional or historical ties. Haitian Creole (ht) overlaps with French, reflecting shared roots Prou et al. (2012). Surprisingly, low-resource languages like Quechua (qu), Haitian Creole (ht), and Swahili (sw) also show overlap, possibly indicating more general or shared neurons.
@@ 6 Conclusion
[…]
@@ A.1 Layer-Wise Amplification on LSS Score
We perform layer-wise amplification on Gemma2 2B model (26 layers) on LSS score using the reconstructed MLAMA dataset. Consistent throughout each language of intervention as captured by the average LSS scores across all languages in Figure 70, the highest is found in 19th layer, and the lowest is in the 10th layer.
Later layers are more likely to result in higher LSS score, especially from the 15th layer until just before the last layer. First few layers may also result in high LSS scores, though not as high. The middle layers result in relatively low LSS scores, often reaching 0 in most languages, with the exception of tr whose the lowest LSS is found in the 18th layer.
This result aligns with findings from Tang et al. (2024) and Kojima et al. (2024) as language-specific neurons are mostly found in the first and last layers, excluding the very last one. Results for layer-wise amplification for each language are provided in Figure 71 (en), 73 (et), 74 (nl), 75 (id), 72 (ms), 76 (vi), 77 (jp), 78 (zh), 79 (fr), 80 (pt), 81 (ru), 82 (it), 83 (ta), 84 (th), 85 (tr).
@@ B.1 Models and Language Neurons Used
We use models trained in different primary languages: Qwen2.5 Qwen et al. (2025) 0.5B and 7B (primarily trained in Chinese and English), Gemma2 Team et al. (2024) 2B and 9B (primarily trained in English), SeaLLMs-v3 Zhang et al. (2024) 1.5B and 7B (primarily trained in Southeast Asian languages). Qwen2.5 and SeaLLMs-v3 share the same model architecture, whereas Gemma2 has a different architecture. We employ neurons specific to 18 languages, English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Haitian Creole (ht), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (th), and Turkish (tr). The selected languages account for different language domains associated with the models and represent both high and low-resource languages to further analyze their behavior.
@@ B.2 Experimental Settings for Finding the Most Optimal Steering Factors
To find the most optimal steering factors, we reconstruct MLAMA dataset Kassner et al. (2021) using 15 languages due to limitation in the dataset. The 15 languages are English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Italian (it), Tamil (ta), Thai (th), and Turkish (tr). In the reconstructed MLAMA dataset, each question has 15 answers in different languages, with no lexical overlap. For example, if both the Indonesian and Malaysian answers were “Jerman” (i.e., Germany in English), that row was discarded.
@@ B.3 Experimental Settings for Task Evaluation
    "Translate from English to Indonesian
@@ Appendix E LSS Scores
The complete LSS scores for both baseline and LAPE neurons are described in Table 5 to Table 18. The aggregated LSS scores for each steering factor is presented in Table 26. Further statistical testing to determine the optimal steering factor is provided in 19. Visualization of token log-probs for same-meaning answers in different languages are provided in both intervened and non-intervened LLMs in Figure 86 and 87.
@@ Appendix F Perplexity Changes After Amplification
Perplexity changes for pmax intervention of LAPE neurons are provided in Figure 60, 61, 58, 59, 62, and 63; pmedian intervention are provided in Figure 66, 65, and 64.
Perplexity changes for pmax intervention of Baseline neurons are provided in Figure 68, 67, and 69.
@@ Appendix G BLEU Score Changes for Translation Tasks After Amplification
For the self-language intervention, BLEU score changes for pmax intervention of LAPE neurons with non-targeted prompt is provided in Table 21. Sample of BLEU delta heatmap (self- and cross-intervention) is provided in Figure 9. BLEU score changes for pmedian intervention of LAPE neurons with targeted prompt is provided in Table 20. pmax intervention on Baseline neurons with targeted prompt is provided in Table 22. Samples of translation output for both targeted and non-targeted prompts are presented in Figure 88 and 89.
@@ Appendix H Delta Accuracies of Reasoning and Knowledge Tasks After Amplification
Table 23 reveals the baseline (no intervention) XWinograd accuracy. Delta accuracies of XWinograd dataset under LAPE neurons when steered with pmax are provided in Figure 21, 22 and 23; intervention results for pmedian are provided in Figure 24, 25, and 26. Delta XWinograd accuracies of pmax intervention on Baseline neurons are provided in Figure 29, 31, and 33.
Table 24 reveals the baseline (no intervention) XCOPA accuracy. Delta accuracies of XCOPA dataset under LAPE neurons when steered with pmax are shown in Figure 34, 35, 36, and 38; intervention results for pmedian are shown in Figure 41, 40, and 42. Delta XCOPA accuracies of pmax intervention on Baseline neurons are provided in Figure 43, 44, and 45.
Table 25 reveals the baseline (no intervention) Include-lite accuracy. Delta accuracies of Include-lite dataset under LAPE neurons when steered with pmax are shown in in Figure 49, 48, 47, 46, 51, and 50, intervention results for pmedian are provided in Figure 53, 52, and 54. Delta Include-lite accuracies of pmax intervention on Baseline neurons are provided in Figure 56, 55, and 57.

### v3

@@ 5.2.3 Perplexity Changes
Self-intervention of LAPE neurons using pmax can modestly reduce perplexity in most languages, showing success in improving language modeling performance as illustrated in Figure 7. Perplexity tends to increase more in cross-intervention settings than in self-intervention. However, the extent of the increase is not always related to the similarity of the language pair; For example, intervention between highly similar and same-family languages like id and ms does not lower perplexity, while less-related pairs like fr on en can. Interestingly, smaller models exhibit larger increases in perplexity, a pattern consistent with findings from the translation task intervention. This may suggest that smaller models contain more localized neurons responsible for language identity (see Appendix F for additional results).
@@ 5.3 Overlap Patterns in Language-Specific Neurons
As illustrated in Figure 8, although LAPE neurons are inherently tailored to individual languages, overlaps still exist across languages, albeit at significantly lower proportions. Compared with Baseline neurons with higher overlap proportions (Appendix D), the stronger specificity of LAPE neurons significantly improve self-intervention and capture patterns in cross-intervention for perplexity and task performance, indicating that better specificity leads to better performance for neuron-level intervention.
A strong overlap is captured in LAPE neurons across mutually intelligible languages like Indonesian (id) and Malaysian (ms), likely due to their linguistic similarity Ahmad and Pustaka (2011). Languages sharing writing systems, such as Chinese (zh) and Japanese (jp) Taylor et al. (1995), also show overlap, consistent with Kojima et al. (2024). Romance languages, i.e., Italian (it), French (fr), and Portuguese (pt), share neurons, and Vietnamese (vi) overlaps moderately with Thai (th) and slightly with id, ms, jp, and zh, likely due to regional or historical ties. Haitian Creole (ht) overlaps with French, reflecting shared roots Prou et al. (2012). Surprisingly, low-resource languages like Quechua (qu), Haitian Creole (ht), and Swahili (sw) also show overlap, possibly indicating more general or shared neurons.
@@ 6 Conclusion
This study proposes the Language Steering Shift (LSS) metric to compare steering factors of language-specific (LAPE) neuron intervention and identify the optimal one, then evaluates its impact in downstream tasks and language modeling. Optimal amplification steers outputs toward target languages with over $90\%$ success on average. Self-language intervention modestly improves or maintains performance in language modeling measured by perplexity, translation, knowledge, and reasoning tasks. Cross-language intervention generally lowers performance, though gains can appear between closely related languages in perplexity and BLEU scores for translation. Compared to less specific language-activated neurons, LAPE neurons yield clearer, more consistent improvements, suggesting that greater neuron specificity enhances multilingual performance when intervened. Overall, these findings suggest that amplifying LAPE neurons in self-language can enhance a model’s language capabilities, although not always improving task accuracy. Furthermore, the consistent degradation from cross-language interventions highlights their limited effectiveness for cross-lingual transfer.
@@ A.1 Layer-Wise Amplification on LSS Score
We perform layer-wise amplification on Gemma2 2B model (26 layers) on LSS score using the reconstructed MLAMA dataset. Consistent throughout each language of intervention as captured by the average LSS scores across all languages in Figure 68, the highest average is found in the 19th layer, and the lowest is in the 10th layer.
Later layers are more likely to result in higher LSS score, especially from the 15th layer until just before the last layer. First few layers may also result in high LSS scores, though not as high. The middle layers result in relatively low LSS scores, often reaching 0 in most languages, with the exception of tr whose the lowest LSS is found in the 18th layer. We observe that some languages from the same family share similar layer-wise LSS scores, such as id, ms, vi and pt, it, fr.
This result aligns with findings from Tang et al. (2024) and Kojima et al. (2024) as language-specific neurons are mostly found in the first and last layers, excluding the very last one. Results for layer-wise amplification for each language are provided in Figure 69 (en), 71 (et), 72 (nl), 73 (id), 70 (ms), 74 (vi), 75 (jp), 76 (zh), 77 (fr), 78 (pt), 79 (ru), 80 (it), 81 (ta), 82 (th), 83 (tr).
@@ B.1 Models and Language Neurons Used
We use models trained in different primary languages: Qwen2.5 Instruct Qwen et al. (2025) 0.5B and 7B (primarily trained in Chinese and English), Gemma2 Instruct Team et al. (2024) 2B and 9B (primarily trained in English), SeaLLMs-v3 Chat Zhang et al. (2024) 1.5B and 7B (primarily trained in Southeast Asian languages). Qwen2.5 and SeaLLMs-v3 share the same model architecture, whereas Gemma2 has a different architecture. We employ neurons specific to 18 languages, English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Haitian Creole (ht), Italian (it), Quechua (qu), Swahili (sw), Tamil (ta), Thai (th), and Turkish (tr). The selected languages account for different language domains associated with the models and represent both high and low-resource languages to further analyze their behavior.
@@ B.2 Experimental Settings for Finding the Most Optimal Steering Factors
To find the most optimal steering factors, we reconstruct MLAMA dataset Kassner et al. (2021) using 15 languages due to limitation in the dataset. The 15 languages are English (en), Dutch (nl), Indonesian (id), Malay (ms), Vietnamese (vi), Japanese (jp), Chinese (zh), French (fr), Portuguese (pt), Russian (ru), Estonian (et), Italian (it), Tamil (ta), Thai (th), and Turkish (tr). In the reconstructed MLAMA dataset, each question has 15 answers in different languages, with no lexical overlap.
@@ B.3 Experimental Settings for Task Evaluation
    "Translate from English to Indonesian.
@@ Appendix E LSS Scores
The complete LSS scores for both Baseline and LAPE neurons are described in Table 5 to Table 18. The aggregated LSS scores for each steering factor is presented in Table 26. Further statistical testing to determine the significance of the optimal steering factor is provided in 19. Visualization of token log-probs for same-meaning answers in different languages are provided in both intervened and non-intervened LLMs in Figure 84 and 85.
@@ Appendix F Perplexity Changes After Amplification
Perplexity changes for pmax intervention of LAPE neurons are provided in Figure 58, 59, 56, 57, 60, and 61; pmedian intervention are provided in Figure 64, 63, and 62. Perplexity changes for pmax intervention of Baseline neurons are provided in Figure 66, 65, and 67.
@@ Appendix G BLEU Score Changes for Translation Tasks After Amplification
For the self-language intervention, BLEU score changes for pmax intervention of LAPE neurons with non-targeted prompt is provided in Table 21. Sample of BLEU delta heatmap (self- and cross-intervention) is provided in Figure 9. BLEU score changes for pmedian intervention of LAPE neurons with targeted prompt is provided in Table 20. pmax intervention on Baseline neurons with targeted prompt is provided in Table 22. Samples of translation output for both targeted and non-targeted prompts are presented in Figure 86 and 87.
@@ Appendix H Delta Accuracies of Reasoning and Knowledge Tasks After Amplification
Table 23 reveals the baseline (no intervention) XWinograd accuracy. Delta accuracies of XWinograd dataset under LAPE neurons when steered with pmax are provided in Figure 21, 22 and 23; intervention results for pmedian are provided in Figure 24, 25, and 26. Delta XWinograd accuracies of pmax intervention on Baseline neurons are provided in Figure 29, 30, and 31.
Table 24 reveals the baseline (no intervention) XCOPA accuracy. Delta accuracies of XCOPA dataset under LAPE neurons when steered with pmax are shown in Figure 32, 33, 34, and 36; intervention results for pmedian are shown in Figure 39, 38, and 40. Delta XCOPA accuracies of pmax intervention on Baseline neurons are provided in Figure 41, 42, and 43.
Table 25 reveals the baseline (no intervention) Include-lite accuracy. Delta accuracies of Include-lite dataset under LAPE neurons when steered with pmax are shown in Figure 47, 46, 45, 44, 49, and 48, intervention results for pmedian are provided in Figure 51, 50, and 52. Delta Include-lite accuracies of pmax intervention on Baseline neurons are provided in Figure 54, 53, and 55.